OpenStudy (anonymous):

Solve the system by substitution. 2x - y + z = -4 z = 5 -2x + 3y - z = -10

OpenStudy (anonymous):

@findme can u help with this one?

OpenStudy (anonymous):

First plug z in your equations. Because it is given.

OpenStudy (radar):

To make it easy on your self, go ahead and substitute for the z as it is a 5 getting. 2x - y =-9 -2x + 3y =-5.

OpenStudy (radar):

Too bad the problem requires substitution as it is ideal to solve using elimination method.

OpenStudy (ajprincess):

z=5 Substitute this in ur first equation 2x-y+z=-4 2x-y+5=-4 2x-y=-4-5 2x-y=-9 Similarly substitute z=5 in ur third equation -2x + 3y - z = -10 -2x+3y-5=-10 -2x+3y=-10+5 -2x+3y=-5 Nw u have two equations 2x-y=-9-(1) -2x+3y=-5-(2) From equation (1) u get y=2x+9 substitute that in equation (2) -2x+3(2x+9)=-5 -2x+6x+27=-5 Solve for x. Then substitute the value of x in y=2x+9 to find y.

OpenStudy (anonymous):

To use the substitution method, solve for x or y . Lets solve for x. Using the equations radar used. (these are your original ones with z plugged in and simplified)

OpenStudy (anonymous):

how do i find x?

OpenStudy (anonymous):

these are the answers i have (–8, 7, 5) (–8, –7, 5) (8, –7, 5) (–8, –7, –5)

OpenStudy (ajprincess):

-2x+6x+27=-5 using this equation can u find x?

OpenStudy (anonymous):

i got -6.2 do i round it to -7?

OpenStudy (ajprincess):

i dnt thnk u will get that answrr when u solve it. can u plz show me your work @algebra2sucks?

OpenStudy (anonymous):

i don't think i did it right cause all i did was add them all up and divide it by -5?

OpenStudy (ajprincess):

-2x+6x=4x -5-27=-32 4x=-32 x=-32/4 =-8 Getting this @algebra2sucks?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

would it b -8, 7,5

OpenStudy (ajprincess):

we know that x=-8 and z=5. to find y let us substitute the value of x in y=2x+9. y=2*-8+9 =-16+9 y=-7

OpenStudy (ajprincess):

Do u understand this @algebra2sucks?

OpenStudy (anonymous):

oh ok so -8, -7, 5

OpenStudy (ajprincess):

yup:)